# Iteration Sums of The Euler’s Totient Function Regarding Powers of Fermat Primes

### v1

@@ 3   $\phi(\phi(\phi(p^{a+1})))$   $\phi(\phi(p^{a}))\cdot\phi(p-1)$
    Notice that 1-1 in the parenthesis multiples $\phi(p-1)$, which nicely cancels out with the $\phi(p-1)$ on the second row, a reason for leaving the $\phi(p-1)$ notation. Then, with some algebraic manipulations:
@@
It is demonstrated that if the statement holds for $k=a$, it would also hold for the next number $k=a+1$, so with the base case, it follows that for any positive integer $k$, Theorem 7 applies. ∎
@@ 5.1 Termination of $\phi^{k}(n)$
In the works of Erd¨os, Granvilie, Pomerance, and Spiro, On the Normal Behavior of the Iterates Of some Arithmetic Functions, they proved some theorems regarding iterations of the Totient function, and stated some open problems[5]. Their study was heavily focused on applying the Totient function $k$ times to a number $n$, such that it is the first time $\phi^{k}(n)=1$, which means the iteration is terminated at the first iteration when the iterated Totient value equals 1. They defined the function $k(n)=k$ as the least number such that $\phi^{k}(n)=1$. Some key results include: If $n=2^{j}$, then,
@@
This could be applied to Lemma 3, so that not only we know the sum would be 2p32p-3, but also we could know that there will be a total of
@@
terms in the sum of Totient values. Another result they found is if $n=2\cdot 3^{j}$, then
@@
where $\lceil x\rceil$ denotes the least integer greater or equal to $x$. This could be applied to Corollary 7.1 to figure out the total number of terms in that series. Furthermore, the case $n=2^{j}$ and $n=2\cdot 3^{j}$ is actually the two extreme behaviors of k(n)k(n)[5], so
@@
for any positive integer $n$. Further investigations could be done to determine the total number of terms in Theorem 7, the series of Totient values on Fermat Primes to the $k$th power.
An Arithmetic Function Arising From the $\phi$ Function, an article from Princeton University by Harold Shapiro, also studied the termination of the iterated Totient function[7]. However, his definition of the last iterate is a little different. He established similar results as Theorem 5 — the iterated Totient value always arrives at 2 — so he defined the function $C(n)=x$ such that $\phi^{x}(n)=2$. One key result they produced with strong relation to this paper is that
@@ 5.2 Sum of Totients
There are also other approaches to summing up Totient functions, and interesting theorems related. For example, in the book History of the Theory of Numbers by Leonard Eugene Dickson[13], a theorem is presented by summing up Totient values from 1 to $n$:
@@
Here an interesting connection arise. Proved by Euler in his Desummis Serierum Reciprocarum, the infinite series $\frac{1}{n^{2}}$ converges to $\frac{\pi^{2}}{6}$[10],
@@
which is the reciprocal of the Totient sums. Although these two topics are far from related, the results show a strong connection between the Totient sums and the $\frac{1}{n^{2}}$ series. Future investigations on the reason behind such a connection could be done, which might uncover some interesting patterns.
@@
This paper proposes closed-form formulas for the sums of iterated Totient functions. To establish context for proving the results, relevant topics were built from the bottom. First, A broad picture of number theory and functions were introduced, which led to the core of this paper, the Euler’s Totient Function. The name and purpose of the function was explained, followed by introducing iteration approaches on the Totient function. Then, the reason behind this paper’s investigations were explained, marking the significance of innovations in number theory. For context and credibility, the historical backgrounds, along with the development of the Totient function and its iteration were presented, summarizing numerous primary sources. Then, definitions were presented. The first definition of the relatively prime relationship between numbers helped the understanding of how the Totient function is defined. The Linear congruence Lemma is first established, to assist in the proof of the Chinese remainder theorem, which then acted as a key function in to prove the multiplicativity of the Totient function. Once it is established that the Totient function is multiplicative, the formula for $\phi(n)$ comes naturally, with two corollaries following. The second corollary was used immediately in the proof of Totient function’s convergence, while the first corollary were used multiple times in the proof of the major theorem in this paper, the iterated Totient sums of powers of Fermat Primes:
$\phi(p^{k})+\phi(\phi(p^{k}))+\phi(\phi(\phi(p^{k})))\cdots+\phi(2)=\frac{2}{p+1}\left[p^{k}(p-1)+2\left(\frac{p-1}{2}\right)^{k}\right]-1.$
This Theorem produces a beautiful corollary for the first Fermat Prime, the number 3:
@@
Overall, the approach of iteration sums opens up a new topic of investigation in number theory. Many theorems and generalizations await to be uncovered.

### v3

@@ 3   $\phi(\phi(\phi(p^{a+1})))$   $\phi(\phi(p^{a}))\cdot\phi(p-1)$
    Notice that 1-1 in the parenthesis multiplies $\phi(p-1)$, which nicely cancels out with the $\phi(p-1)$ on the second row, a reason for leaving the $\phi(p-1)$ notation. Then, with some algebraic manipulations:
@@
We demonstrated that if the statement holds for $k=a$, it would also hold for the next number $k=a+1$, so with the base case, it follows that for any positive integer $k$, Theorem 7 applies. ∎
@@ 5.1 Termination of $\phi^{k}(n)$
In the works of Erdös, Granvilie, Pomerance, and Spiro, On the Normal Behavior of the Iterates Of some Arithmetic Functions, they proved some theorems regarding iterations of the totient function, and stated some open problems [termination_of_phi^k]. Their study was heavily focused on applying the totient function $k$ times to a number $n$, such that it is the first time $\phi^{k}(n)=1$, which means the iteration is terminated at the first iteration when the iterated totient value equals 1. They defined the function $k(n)=k$ as the least number such that $\phi^{k}(n)=1$. Some key results include: If $n=2^{j}$, then,
@@
We could apply this to Lemma 3, so that not only we know the sum would be 2p32p-3, but also we could know that there will be a total of
@@
terms in the sum of totient values series. Another result they found is if $n=2\cdot 3^{j}$, then
@@
where $\lceil x\rceil$ denotes the least integer greater or equal to $x$. We could apply this to Corollary 7.1 to figure out the total number of terms in that series. Furthermore, the case $n=2^{j}$ and $n=2\cdot 3^{j}$ is actually the two extreme behaviors of $k(n)$ [termination_of_phi^k], so
@@
for any positive integer $n$. Further investigations could determine the total number of terms in Theorem 7, the series of totient values on Fermat Primes to the $k$th power.
An Arithmetic Function Arising From the $\phi$ Function, an article from Princeton University by Harold Shapiro, also studied the termination of the iterated totient function [C(n)_function]. However, his definition of the last iterate is a little different. He established similar results as Theorem 5 — the iterated totient value always arrives at 2 — so he defined the function $C(n)=x$ such that $\phi^{x}(n)=2$. One key result they produced with strong relation to this paper is that
@@ 5.2 Sum of Totients
There are also other approaches to summing up totient functions, and interesting theorems related. For example, in the book History of the Theory of Numbers by Leonard Eugene Dickson [History_of_the_theory_of_numbers_6/pi^2], they present a theorem by summing up totient values from 1 to $n$:
@@
Here an interesting connection arise. Proved by Euler in his Desummis Serierum Reciprocarum, the infinite series $\frac{1}{n^{2}}$ converges to $\frac{\pi^{2}}{6}$ [series_1/n^2_pi^2/6],
@@
which is the reciprocal of the totient sums. Although these two topics are far from related, the results show a strong connection between the totient sums and the $\frac{1}{n^{2}}$ series. Future investigations could try to uncover the reason behind such interesting connection
@@ 6 Conclusion
This paper proposes closed-form formulas for the sums of iterated totient functions. First, A broad picture of number theory and functions were introduced, leading to the Euler’s totient Function. We explained the name and purpose of the function, introducing iteration approaches on the totient function. Next, we presented definitions and fundamental theorems including relatively prime, multiplicative functions, the Linear congruence Lemma, and the Chinese remainder theorem. We prove the formula to calculate $\phi(n)$, with two corollaries following. We used the second corollary immediately in the proof of totient function’s convergence, while we used the first corollary multiple times in the proof of the iterated totient sums of Fermat Primes, the main result of this paper:
$\phi(p^{k})+\phi(\phi(p^{k}))+\phi(\phi(\phi(p^{k})))+\cdots+\phi(2)=\frac{2}{p+1}\left[p^{k}(p-1)+2\left(\frac{p-1}{2}\right)^{k}\right]-1.$
This theorem produces a clean corollary for the first Fermat Prime, the number 3:
@@
Overall, the approach of iteration sums opens up a new topic to investigate. Many theorems and generalizations await to be uncovered.